# Score Distillation via Reparametrized DDIM

### v1

@@ 7 Conclusion, Limitations, and Future Work
Some limitations of our algorithm motivate future work. While we improve the sample quality of each view, 3D consistency between views remains challenging; as a result, despite the convexity loss, our algorithm occasionally produces flat or concave “billboards.” A possible resolution might involve supervision with pre-trained depth or normal estimators. A related problem involves content drift from one view to another. Since there was no 3D supervision, there is little to no communication between opposite views, which can lead to inconsistent 3D assets. Stronger view conditioning, multi-view supervision, or video generation models might resolve this problem. Finally, score distillation is capped by the performance of the underlying diffusion model and is hence prone to reproduce similar “hallucinations” (e.g., text and limbs anomalies); the algorithm inherits the biases of the 2D diffusion model and can produce skewed distributions. Section E.1 demonstrates typical failure cases.
@@ Acknowledgments
         $\epsilon\leftarrow\kappa^{t+\tau}(g(\psi,c))$
@@
Our implementation uses the following choice of $\kappa^{t}$:
$\begin{split}H(t)&=0.3\sqrt{1-\alpha(t)}\epsilon_{H}\text{~{}, where~{}}\epsilon_{H}\sim\mathcal{N}(0,I)\\ \kappa^{t}(x)&=\textrm{ddim\_inversion}(x,t)+H(t)\end{split}$ (11)
@@
Section E.1 demonstrates examples of typical failure cases, including the Janus problem and the hollow face illusion.
@@
We implement our algorithm in threestudio [threestudio2023] on top of SDS [poole2022dreamfusion]. We use Stable Diffusion 2.1 [Rombach_2022_CVPR] as the diffusion model. For volumetric representation we use InstantNGP [muller2022instant]. Instead of randomly sampling time $t$ as in SDS, we maintain a global parameter $t$ that linearly decays from $1$ to $0.2$ (lower time values do not have a significant contribution). Next, for each step we render a $512\times 512$ random view and infer $\kappa^{t+\tau}$ by running DDIM inversion for $\texttt{int}(10t)$ steps. We use NVIDIA A6000 GPUs and run each generation for $10k$ steps with learning rate of $10^{-2}$, which takes approximately 2 wall-clock hours per shape generation.
@@ A.4 Prompts used in the quantitative evaluation
Appendix B CFG for inversion in 2D
@@
Figure 13 shows that when inverting the DDIM trajectory starting with a clean image, the best inversion strategy uses $\gamma_{\text{inv}}=1$ and then regenerates the image via DDIM with $\gamma_{\text{fwd}}=1$. [miyake2023negative] suggests this approach, which inverts the computation almost perfectly. Other approaches introduce bias and only approximate the image on the forward pass.
As we have seen in fig. 9, in 3D, the story is different: $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=1$ fails to generate any meaningful shape. We explain this behavior by the fact that in our 2D experiment, the original image contains class information; inverting with $\gamma_{\text{inv}}=1$ encodes this information into the noise and thus does not need additional CFG during reconstruction. In 3D generation, however, the initial NeRF does not contain class information and needs bigger CFG on the forward pass. Additionally, fig. 13 demonstrates the inversion quality of an entire DDIM trajectory, whereas eq. 10 we are interested in inverting a single-step denoising only. what might explain the fact that seemingly bad inversion quality of $\gamma_{\text{fwd}}=-\gamma_{\text{inv}}=7.5$ yields much better results in 3D.
@@ Appendix C ODE derivation
Appendix D Additional Intuition: Out-of-Distribution Correction
@@
Appendix E Additional Results
E.1 Failure Cases
We provide illustrations of typical failure cases of our algorithm in Figures 15, 16, 17 and 18.
@@ E.1 Failure Cases
E.2 Additional generations
Figures 19, 20, 21 and 22 provide additional generations produced using our method.
@@ E.2 Additional generations
E.3 Additional comparison with baselines
We provide additional qualitative comparisons with other score distillation algorithms in figs. 23, 24 and 25.

### v2

@@ 7 Conclusion, Limitations, and Future Work
Some limitations of our algorithm motivate future work. While we improve the sample quality of each view, 3D consistency between views remains challenging; as a result, despite the convexity loss, our algorithm occasionally produces flat or concave “billboards.” A possible resolution might involve supervision with pre-trained depth or normal estimators. A related problem involves content drift from one view to another. Since there was no 3D supervision, there is little to no communication between opposite views, which can lead to inconsistent 3D assets. Stronger view conditioning, multi-view supervision, or video generation models might resolve this problem. Finally, score distillation is capped by the performance of the underlying diffusion model and is hence prone to reproduce similar “hallucinations” (e.g., text and limbs anomalies); the algorithm inherits the biases of the 2D diffusion model and can produce skewed distributions. Section F.1 demonstrates typical failure cases.
@@ Acknowledgments
         $\epsilon\leftarrow\kappa^{t+\tau}_{y}(g(\psi,c))$
@@
Our implementation uses the following choice of $\kappa^{t}_{y}$:
$\begin{split}H(t)&=0.3\sqrt{1-\alpha(t)}\epsilon_{H}\text{~{}, where~{}}\epsilon_{H}\sim\mathcal{N}(0,I)\\ \kappa^{t}_{y}(x)&=\textrm{ddim\_inversion}(x,t)+H(t)\end{split}$ (11)
@@
Section F.1 demonstrates examples of typical failure cases, including the Janus problem and the hollow face illusion.
@@
We implement our algorithm in threestudio [threestudio2023] on top of SDS [poole2022dreamfusion]. We use Stable Diffusion 2.1 [Rombach_2022_CVPR] as the diffusion model. For volumetric representation we use InstantNGP [muller2022instant]. Instead of randomly sampling time $t$ as in SDS, we maintain a global parameter $t$ that linearly decays from $1$ to $0.2$ (lower time values do not have a significant contribution). Next, for each step we render a $512\times 512$ random view and infer $\kappa^{t+\tau}$ by running DDIM inversion for $\texttt{int}(10t)$ steps - i.e. we use 10 steps of DDIM inversion for $t=1$ and linearly decrease it for smaller $t$. We use NVIDIA A6000 GPUs and run each generation for $10k$ steps with learning rate of $10^{-2}$, which takes approximately 2 wall-clock hours per shape generation.
@@ A.4 Prompts used in the quantitative evaluation
Appendix B Visual comparison of noise patterns
In this section, we provide additional examples of the noise samples obtained with different $\kappa_{y}^{t}$. As can be seen from fig. 13, optimization with SGD almost does not change the noise pattern compared to random sampling. Fixed-point iterations drift in a notable but counterproductive direction, and the single-step prediction from the obtained noisy images (depicted on the right side) degrades. DDIM inversion with $\gamma_{\text{inv}}=7.5$ and $\gamma_{\text{inv}}=-7.5$ produces similar noise patterns located around the “meaningful” parts of the image; however, $\gamma_{\text{inv}}=-7.5$ results in noise patterns that are much more compatible with the single-step denoising procedure used in SDS, leading to much more accurate details appearing for earlier $t$.
Figure 13: Examples of noisy images (left part) obtained for different choices of $\kappa_{y}^{t}$. On the right part, a single step denoised images are presented for the corresponding noisy images.
Appendix C CFG for inversion in 2D
@@
Figure 14 shows that when inverting the DDIM trajectory starting with a clean image, the best inversion strategy uses $\gamma_{\text{inv}}=0$ and then regenerates the image via DDIM with $\gamma_{\text{fwd}}=0$. [miyake2023negative] suggests this approach, which inverts the computation almost perfectly. Other approaches introduce bias and only approximate the image on the forward pass.
As we have seen in fig. 9, in 3D, the story is different: $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=0$ fails to generate any meaningful shape. We explain this behavior by the fact that in our 2D experiment, the original image contains class information; inverting with $\gamma_{\text{inv}}=0$ encodes this information into the noise and thus does not need additional CFG during reconstruction. In 3D generation, however, the initial NeRF does not contain class information and needs bigger CFG on the forward pass. Additionally, fig. 14 demonstrates the inversion quality of an entire DDIM trajectory, whereas eq. 10 we are interested in inverting a single-step denoising only. what might explain the fact that seemingly bad inversion quality of $\gamma_{\text{fwd}}=-\gamma_{\text{inv}}=7.5$ yields much better results in 3D.
@@ Appendix D ODE derivation
Appendix E Additional Intuition: Out-of-Distribution Correction
@@
Appendix F Additional Results
F.1 Failure Cases
We provide illustrations of typical failure cases of our algorithm in Figures 16, 17, 18 and 19.
@@ F.1 Failure Cases
F.2 Additional generations
Figures 20, 21, 22 and 23 provide additional generations produced using our method.
@@ F.2 Additional generations
F.3 Additional comparison with baselines
We provide additional qualitative comparisons with other score distillation algorithms in figs. 24, 25 and 26. We compare with Dreamfusion [poole2022dreamfusion], Magic3D [lin2023magic3d], Fantasia3D [chen2023fantasia3d], Stable Score Distillation (StableSD [tang2023stable], Classifier Score Distillation (ClassifierSD [yu2023text]), Noise-Free Score Distillation (NFSD [katzir2023noise]), ProlificDreamer or Variational Score Distillation (VSD [wang2024prolificdreamer]), Interval Score Matching (ISM) [liang2023luciddreamer], and HiFA [zhu2023hifa].